OpenStudy (anonymous):

Please help me? Sharon earns $25 per item she sells plus a base salary of $100 per week. Write and solve an inequality to find how many items she must sell to earn at least $700 per week. Provide your conclusion as a complete sentence.

OpenStudy (jack1):

if she was to earn exactly $700 per week = base salary $100 + 25(x) where x is the number of items she sells in a week 700 = 100 +25x 600 = 25x so x = 600/25 =? your inequality eqn can be gleaned from there, yes?